Gauge no. 36 copper wire has a diameter of 0.058m. Find the length of no. 36 copper wire which will give a resistance of 0.0005 ohms


1
Expert's answer
2021-05-25T10:10:26-0400

"R=\rho l/A"


"A=\pi d^2/4=3.14\cdot0.058^2/4=0.00264\ (m^2)"


"l=R\cdot A/\rho=0.0005\cdot 0.00264/1.75\cdot10^{-8}=75.45\ (m)" . Answer